OpenStudy (anonymous):

factor 2x^2-x=3 could someone please explain with steps I'm really bad at factoring

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

WJHAT

OpenStudy (anonymous):

SOMEONE HELP ME PLEASE

OpenStudy (jonnyvonny):

Dang

OpenStudy (jonnyvonny):

I typed out this long, elaborate way to do this, and I accidently clicked on another person's question qq

OpenStudy (jonnyvonny):

Well, bear me for a moment again, please.

OpenStudy (anonymous):

LOL TY ANYWAYS

OpenStudy (ranga):

2x^2-x=3 2x^2 - x - 3 = 0 2x^2 - 3x + 2x - 3 = 0 x(2x - 3) + 1(2x - 3) = 0 (2x - 3)(x + 1) = 0

OpenStudy (jonnyvonny):

First off, you must set the equation equal to 0:\[2x^2-x-3=0\] From then, upon seeing the -3, you know that the factors of 3 (which are 1 and 3), upon multiplying them, must be negative, and the only way to get a negative value from multiplying, is for one of the values to be positive and the other negative. Secondly, since you see the -1 (-x has an invisible -1 in front of it), you must conclude that the negative value must be greater. However, you must keep the "2" in front of the x^2 in mind. Now, as we said before, the values of 3 are 1 and 3, you may think, what value, multiplied by 2 makes -1? We see that -3+2=-1;\[(2x-3)(x+1)\] If you're not too sure about the answer, distribute. I hope this helped, pst me if you need more assistance with this or any other problem. :)

OpenStudy (anonymous):

my answer key's answer that is closest to that is 1,3 is this correct

OpenStudy (anonymous):

@JonnyVonny

OpenStudy (jonnyvonny):

Mmk, strange, because if you plug in 1 into the original equation, it does not equal 0, which it is suppose to.